OpenStudy (anonymous):

The length of a water color painting is 5 inches less than twice its width.If the perimeter of the painting is 62 inches,what is its width in inches? Choose from one of these:10,8,12,none of the above

OpenStudy (anonymous):

Will need help

OpenStudy (anonymous):

Can someone help me

OpenStudy (raden):

l = 2w-5 ... (1) 2l + 2w = 62, it can simplified by divided by 2 to both sides, gives l + w = 31 .... (2) now, substitute 1st equation to the 2nd, what u get

OpenStudy (anonymous):

Now this is a ruff one for me to figure out

OpenStudy (anonymous):

Still need help

OpenStudy (raden):

well, if u use subtitution method (1) subs to (2), gives 2w-5+w=31 3w=36 w=12

OpenStudy (kainui):

First off, you should consider approaching this problem differently. Think of everything you're doing as simplifying it down. So I'll help walk you through this. We'll start with the first sentence. "The length of a water color painting is 5 inches less than twice its width." So let's see, what can we simplify down here? When they say the length, can we not say "L" to make it smaller and easier to look at? \[L\] Ok, what's next? They say the length _is_ something. Well, there's your hint that length (L) is equal to something, so let's update our formula... \[L=\] Ok, now what's it say? 5 inches less than twice the width. Ok. So we can do the same thing for width here as we did for length by calling it W. \[L=W????\] 5 inches less than twice the width. We know less than is subtracting and twice is multiplying times two, right? So we can just put in a 2 next to the W to show we doubled the width and -5 to show that it's less than that. \[L=2W-5\] Now try it on your own for the second part.